OpenStudy (anonymous):

I need an URGENT answer PLEASE! Rukia mixed 2 holds 5/18 litres of blue paint with 1 holds 5/15 litres of white paint to make a sky blue paint. How many litres of sky blue paint did she make?

OpenStudy (anonymous):

8/9 i think....not sure what you mean by "holds" though

OpenStudy (anonymous):

2 holds 5/18 litres of blue paint = 2*(5/18) litres = 10/18 litres = 5/9 litres 1 holds 5/15 litres of white paint = 1*(5/15) = 5/15 litres = 1/3 litres so sky blue paint = blue paint + white paint = 5/9 + 1/3 simplify this by taking LCM (LCD) and we get (5/9) + (1/3)*(3/3) =5/9 + 3/9 =5+3/9 =8/9